# Teaching LLMs to Abstain across Languages via Multilingual Feedback

### v1

@@ Abstract
Multilingual large language models (LLMs) often have knowledge disparities across languages, with larger gaps in under-resourced languages. Teaching LLMs to abstain in the face of knowledge gaps is thus a promising strategy to mitigate hallucinations in multilingual settings. However, previous studies on LLM abstention primarily focus on English; we find that directly applying these solutions beyond English results in up to 20.5% performance gaps between high and low-resource languages, potentially due to LLMs’ drop in calibration and reasoning beyond a few resource-rich languages. To this end, we propose strategies to enhance LLM abstention by learning from multilingual feedback, where LLMs self-reflect on proposed answers in one language by generating multiple feedback items in related languages: we show that this helps identify the knowledge gaps across diverse languages, cultures, and communities. Extensive experiments demonstrate that our multilingual feedback approach outperforms various strong baselines, achieving up to 9.2% improvement for low-resource languages across three black-box and open models on three datasets, featuring open-book, closed-book, and commonsense QA. Further analysis reveals that multilingual feedback is both an effective and a more equitable abstain strategy to serve diverse language speakers, and cultural factors have great impact on language selection and LLM abstention behavior, highlighting future directions for multilingual and multi-cultural reliable language modeling. [^1] Code and data will be publicly available at https://github.com/BunsenFeng/M-AbstainQA.
@@
Shangbin Feng1   Weijia Shi1   Yike Wang2   Wenxuan Ding3   Orevaoghene Ahia1 Shuyue Stella Li1    Vidhisha Balachandran4    Sunayana Sitaram5    Yulia Tsvetkov1 1University of Washington   2University of California, Berkeley 3The Hong Kong University of Science and Technology 4Carnegie Mellon University   5Microsoft Research shangbin@cs.washington.edu
@@ 1 Introduction
Figure 1: Average accuracy of abstention baselines in low- and high-resource languages with Aya-13B. Existing abstain strategies drop by up to 20.5% when applied to QA in low-resource languages.
@@ 7 Conclusion
When retrieval corpora are not readily available in low-resource languages, cross-lingual retrieval (Asai et al., 2021; Shen et al., 2022; Huang et al., 2023c; Wieting et al., 2023; Lin et al., 2023; Thakur et al., 2024) is often necessary for retrieval-augmented LLMs, where user queries are translated to high-resource languages and retrieval is performed with that language. We investigate whether our multilingual feedback approach works in this setting: we use English Wikipedia for retrieval  [^5] We retrieve Wikipedia with the WikiSearch API. and prepend back-translated paragraphs before the query from the seven low-resource languages. We evaluate various abstain approaches with ChatGPT and present performance in Figure 7. Our proposed multilingual feedback approach outperforms baselines for six of the seven low-resource languages, by 6.9% on average. This indicates that our Multi-related approach could also improve multilingual LLM reliability in retrieval-augmented settings.
@@ Appendix B Experiment Details
We present the related languages employed for feedback generation in the Language Relatedness study (§5) in Tables 10, 11, and 12.

### v3

@@ Abstract
Multilingual large language models (LLMs) often have knowledge disparities across languages, with larger gaps in under-resourced languages. Teaching LLMs to abstain in the face of knowledge gaps is thus a promising strategy to mitigate hallucinations in multilingual settings. However, previous studies on LLM abstention primarily focus on English; we find that directly applying these solutions beyond English results in up to 20.5% performance gaps between high and low-resource languages, potentially due to LLMs’ drop in calibration and reasoning beyond a few resource-rich languages. To this end, we propose strategies to enhance LLM abstention by learning from multilingual feedback, where LLMs self-reflect on proposed answers in one language by generating multiple feedback items in related languages: we show that this helps identify the knowledge gaps across diverse languages, cultures, and communities. Extensive experiments demonstrate that our multilingual feedback approach outperforms various strong baselines, achieving up to 9.2% improvement for low-resource languages across three black-box and open models on three datasets, featuring open-book, closed-book, and commonsense QA. Further analysis reveals that multilingual feedback is both an effective and a more equitable abstain strategy to serve diverse language speakers, and cultural factors have great impact on language selection and LLM abstention behavior, highlighting future directions for multilingual and multi-cultural reliable language modeling. [^1] Code and data are publicly available at https://github.com/BunsenFeng/M-AbstainQA.
@@
Shangbin Feng1   Weijia Shi1   Yike Wang1   Wenxuan Ding2   Orevaoghene Ahia1 Shuyue Stella Li1    Vidhisha Balachandran3    Sunayana Sitaram3    Yulia Tsvetkov1 1University of Washington   2The University of Texas at Austin   3Microsoft Research shangbin@cs.washington.edu
@@ 1 Introduction
Figure 1: Average accuracy of abstention baselines in low- and high-resource languages with Aya-13B. Existing abstain strategies drop by 8.4% on average when applied to QA in low-resource languages.
@@ 7 Conclusion
Acknowledgements
We gratefully acknowledge support from the National Science Foundation under CAREER Grant No. IIS2142739, and NSF grants No. IIS2125201 and IIS2203097. This work was also supported in part by gift funding from Google, MSR, and Microsoft Accelerate Foundation Models Research.
@@ Acknowledgements
When retrieval corpora are not readily available in low-resource languages, cross-lingual retrieval (Asai et al., 2021; Shen et al., 2022; Huang et al., 2023c; Wieting et al., 2023; Lin et al., 2023; Thakur et al., 2024) is often necessary for retrieval-augmented LLMs (Lewis et al., 2020; Shi et al., 2024; Yasunaga et al., 2023; Xu et al., 2024), where user queries are translated to high-resource languages and retrieval is performed with that language. We investigate whether our multilingual feedback approach works in this setting: we use English Wikipedia for retrieval  [^5] We retrieve Wikipedia with the WikiSearch API. and prepend back-translated paragraphs before the query from the seven low-resource languages. We evaluate various abstain approaches with ChatGPT and present performance in Figure 7. Our proposed multilingual feedback approach outperforms baselines for six of the seven low-resource languages, by 6.9% on average. This indicates that our Multi-related approach could also improve multilingual LLM reliability in retrieval-augmented settings.
@@
Randomness in Sampling Feedback
We randomly sample feedback sets with temperature of 1 and repeat for 3 runs. If the LLM abstains/answers in all 3 runs, then it is deemed consistent; 1:2 and 2:1 scenarios are then deemed as inconsistent. We present results in Table 10, showing that learning to abstain from multilingual feedback is largely consistent.
bn ml mr ne ta te
MMLU, consistent 103 101 99 117 114 95
MMLU, inconsistent 42 26 35 31 21 34
consistent rate % 71.03 79.53 73.88 79.05 84.44 73.64
Hellaswag, consistent 115 136 115 131 122 122
Hellaswag, inconsistent 39 14 39 24 18 25
consistent rate % 74.68 90.67 74.68 84.52 87.14 82.99
Table 10: Consistency when repeating the feedback sampling for three times.
@@ Appendix B Experiment Details
We present the related languages employed for feedback generation in the Language Relatedness study (§5) in Tables 11, 12, and 13.